OpenStudy (anonymous):

simplify 3/x+13/x−11 and explain please !

jimthompson5910 (jim_thompson5910):

the problem is \[\large \frac{3}{x} + \frac{13}{x-1}\] right?

OpenStudy (anonymous):

no, its an 11 not a 1 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

oh so \[\large \frac{3}{x} + \frac{13}{x-11}\] sry misread

OpenStudy (anonymous):

yessir @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

The idea is to get each fraction to have the same denominator (the LCD). In this case, the LCD is x(x-11) To get the first fraction have a denominator of x(x-11), you need to multiply top and bottom by (x-11), which is the missing piece here. To get the second fraction have a denominator of x(x-11), you need to multiply top and bottom by x, which is the missing piece here. Afterwards, you can combine the numerators over the common denominator. \[\large \frac{3}{x} + \frac{13}{x-11}\] \[\large \frac{3(x-11)}{x(x-11)} + \frac{13}{x-11}\] \[\large \frac{3x-33}{x(x-11)} + \frac{13}{x-11}\] \[\large \frac{3x-33}{x(x-11)} + \frac{x*13}{x(x-11)}\] \[\large \frac{3x-33}{x(x-11)} + \frac{13x}{x(x-11)}\] \[\large \frac{3x-33+13x}{x(x-11)}\] \[\large \frac{16x-33}{x(x-11)}\] \[\large \frac{16x-33}{x^2 - 11x}\] Note: the last step is (usually) optional

OpenStudy (anonymous):

so "x-11' is the LCD @jim_thompson5910